Compare the values in Column A and Column B for the solution of the system:
{ x + y + z = 10
x + y + 2z = 13
x + y + 3z = 16
Column A = x
Column B = y

A The value in Column A is greater
B The value in Column B is greater
C The two values are equal
D The relationship cannot be determined on the basis of the information supplied
I'm just not sure!!! Thanks!!

Answer by AnlytcPhil(1810)

Compare the values in Column A and Column B for the solution of the system: 

{ x + y + z = 10
x + y + 2z = 13
x + y + 3z = 16
Column A = x
Column B = y
A The value in Column A is greater
B The value in Column B is greater
C The two values are equal
D The relationship cannot be determined on the basis of the information supplied
I'm just not sure!!! Thanks!!
Subtracting the first equation from the second gives z = 3
Substituting z = 3 makes all three equations become x + y = 7
So there is no way to tell whether x or y is larger or if they are the same.
For instance, x could be 6 and y 1 or vice-versa or they could both be 3.5
The correct choice is D
Edwin